OpenStudy (softballgirl372015):

how would you find the range of this function? g(x)=8-2x-x^2

OpenStudy (campbell_st):

well its a parabola thats concave down. so I'd start by finding the line of symmetry. given a parabola \[y = ax^2 + bx + c\] the line of symmetry is \[x = \frac{-b}{2a}\] so in your question \[x = \frac{2}{-2}\] line of symmetry is x = -1 the maximum value the curve gets to is on the line of symmetry. So substitute x = -1 into the equation y = 8 -2(-1) - (-1)^2 so y = 9 the vertex of the parabola is (-1, 9) the maximum value the curve gets to is y = 9 so the range is all values of y less than or equal to 9

OpenStudy (softballgirl372015):

omg thank you sooo much!!!!